# Near-extremal hydrodynamics and the holographic product formula

### v1

@@ 4 An application to holographic neutrino transport
In this last section, the extremal hydrodynamic product formula (19) is applied to the setup considered in Jarvinen:2023xrx. There, a toy model was introduced to investigate neutrino transport in holography, where the charged current spectral functions were computed from the Maxwell equations on the RN background, i.e. the problem considered in section 3.1. The numerical data for the spectral functions was then used to compute the neutrino and anti-neutrino opacities, $\kappa$ and $\bar{\kappa}$. An interesting observation from this calculation was that the opacities at low temperatures were quite well approximated by substituting for the current correlators the leading order hydrodynamic expression Kovtun:2012rj, extrapolated to zero temperature.
@@
Figure 3: Neutrino opacity $\kappa$ and its hydrodynamic approximations as a function of the neutrino energy $E_{\nu}$, for $\mu/T\simeq 4.7$ (left) and $\mu/T\simeq 65$ (right). $\kappa_{e,0}$ is a normalization factor that was introduced in Jarvinen:2023xrx.
@@
The better description of the correlators at low energies translates into a better approximation to the neutrino opacities, as illustrated in figures 3 and 4. The latter show a comparison of the exact holographic opacities with the hydrodynamic and extremal hydrodynamic approximations, as a function of neutrino energy $E_{\nu}$ and for two different values of the ratio of chemical potential to temperature [^17] For $T=10\,\text{MeV}$, these values correspond to baryon number densities $n_{B,1}=10^{-3\,}\mathrm{fm^{-3}}$ and $n_{B,2}=1\,\mathrm{fm^{-3}}$, which covers the typical range expected in neutron stars Jarvinen:2023xrx.: $\mu_{1}/T\simeq 4.7$ and $\mu_{2}/T\simeq 65$. For most values of $E_{\nu}$, the extremal and standard hydrodynamic approximations are observed to have similar accuracies, except near the dip in the neutrino opacity around $E_{\nu}\sim\mu_{\nu}$ (figure 3), where the extremal hydrodynamic approximation gives a significantly better description. This is consistent with expectations, since the integrand from which the neutrino opacity is computed (see Jarvinen:2023xrx) is concentrated near $\omega=0$ in this case. Note that this regime of neutrino energies is particularly important to describe well since it corresponds to the effective Fermi surface [^18] See Jarvinen:2023xrx for a discussion of the quasi-particle approximation considered for the description of neutrinos., near which most neutrinos are emitted and absorbed.
@@ 5 Discussion
In this work, we used the holographic product formula introduced in Dodelson:2023vrw to get insight into the structure of holographic spectral functions in the near-extremal hydrodynamic regime, $\omega,k,T\ll\mu$. In the limit of vanishing temperature $T\ll\omega,k\ll\mu$, the formula provides an expression where both the low temperature gapless modes and the IR scaling behavior appear explicitly, moreover in a simple factorized form. As we showed in sections 3 and 4, the product formula applies in many holographic settings, where it allows to extend the near-extremal hydrodynamic approximation of correlators to the regime of low energies.
@@
  • Another interesting point is related to the recent results of Grozdanov:2025ner; Grozdanov:2025ulc in four bulk dimensions, which showed that duality relations between different parity sectors of correlators translate into additional constraints on their structure. These may allow to write even more useful expressions in the near-extremal hydrodynamic regime of $(2+1)$-dimensional holographic theories;
  • In Dodelson:2023vrw, the UV limit $\omega\gg T$ of the product formula was analyzed in relation with the OPE in the UV CFT. Here, we were rather interested in an IR regime of the correlators $\omega\ll\mu$, but also considered the limit $T\ll\omega\ll\mu$ to write the extremal hydrodynamic product formula. In this case, the corrections of order $\mathcal{O}(T/\omega)$ to the extremal result may still be controlled by an OPE, although rather associated with the IR CFT, which could be worth investigating;
@@
Finally, the main problem remains to understand what is the structure of the low energy effective field theory (EFT) that underlies the near-extremal spectral functions, which should combine both hydrodynamic-like features and the IR CFT. The formalism of Crossley:2015evo; Glorioso:2017fpd; Liu:2018kfw should a priori encompass this kind of theories, since it is well adapted to describe systems with hydrodynamic-like features Blake:2017ris. This type of effective theories are also known to arise in holographic calculations Nickel:2010pr; Glorioso:2018mmw; Davison:2022vqh, which however still remain to be extended to the relevant systems.
@@ Appendix A IR limit of near-extremal correlators
Now, the procedure to exhibit the relation with the IR correlator is based on the standard division of the bulk into inner and outer region Faulkner:2009wj; DHoker:2010xwl:
@@
    where the coefficient of the subleading term is the IR correlator by definition Faulkner:2009wj, and the dots go to zero as $\zeta\to 0$. Note that $\Delta_{\text{IR}}=1/2+\nu$ is (half) the scaling dimension of the IR correlator associated with $\psi$.
  • The outer region is where $\zeta\ll 1$. In particular, it overlaps with the inner region near the boundary of the IR AdSp, where the outer solution behaves as in (127). By linearity of (121), the full outer solution may then be written as Faulkner:2009wj
@@ Appendix B An example at next-to-leading order in the near-extremal hydrodynamic expansion
We discuss here how the correlator computed in Gouteraux:2025kta at next-to-leading order in the near-extremal hydrodynamic expansion fits with the product formula (14), and in particular with its extremal limit (19).
@@
the result of Gouteraux:2025kta takes the form [^22] Note that we also need $i\omega/(2\pi T)\notin\mathbb{N}^{*}$ for $\mathcal{G}$ in (135) to be finite. In other words, the expansion (134) breaks down when $i\omega/(2\pi T)$ is a positive integer.
@@
We are interested in the limit [^23] Note that, in this case, the soft poles do not match the poles of the IR CFT correlator as in (20) Gouteraux:2025kta. There is therefore no simple expression of the form (23) for the correlator in the general near-extremal regime $|\omega|\sim|\vec{k}^{2}|\sim T\ll r_{e}^{-1}$. However, the extremal formula (19) still applies. $\omega/T\to\infty$, for which the product formula (14) simplifies according to (18), whereas (134)-(135) becomes (for $\omega$ off the negative imaginary axis)
@@
At this order equation (136) has two poles Gouteraux:2025kta
@@
with the denominator given by
@@
with the IR scaling dimension given in this case by Gouteraux:2025kta
@@ Appendix C Correlators for different field variables in the probe longitudinal gauge sector
For $d\geq 7$, the leading term $\phi_{(0)}$ is non-normalizable and should be indentified as the source, so that $G_{\Psi}^{R}$ is unambiguously identified as $G_{\Psi}^{R}=\phi_{(\Delta)}/\phi_{(0)}$. However, if $d\leq 6$, the dimension $\Delta$ is inferior to $(d+2)/2$, for which two possible quantizations are possible, depending on whether $\phi_{(0)}$ or $\phi_{(\Delta)}$ is treated as the source Klebanov:1999tb. In this case, two zero-free product formulae may therefore be written: one for the two-sided correlator associated with $\phi_{(\Delta)}/\phi_{(0)}$, and the other for the inverse $\phi_{(0)}/\phi_{(\Delta)}$ Klebanov:1999tb. However, it is not difficult to see that one of these formulae automatically implies the other.
@@ Appendix D Poles of the probe current correlators
In this appendix, we present a numerical calculation of the low-lying (gapless and soft) poles of the RN probe current correlators discussed in section 3.1, in the near-extremal regime. We used for this the Mathematica package of Jansen:2017oag. The transverse and longitudinal sectors are discussed separately.
@@ D.2 Longitudinal correlator
We now present the numerical results for the poles of the longitudinal correlator, which are shown in figures 6 (imaginary part) and 7 (real part), still with $d=4$ and $r_{e}T=(50\pi)^{-1}$. This sector contains both a diffusive pole $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ and a sequence of soft poles, which interact non-trivially. The behavior as momentum increases is very similar to Gouteraux:2025kta, with the diffusive pole splitting into two poles with finite real parts before recombining on the imaginary axis several times as it collides with soft poles, before eventually remaining outside the imaginary axis, with a behavior of the form (138). Figure 6 indicates that the soft poles asymptote to the IR poles (153) (orange lines) at large momenta, but deviate significantly around the point where they cross the diffusive mode.
@@ F.1 $n=2$
The helicity-0 correlators on the AdS4-RN background were studied in Edalati:2010pn. As for $n=3$, the system of fluctuation equations can be decoupled by considering the Kodama-Ishibashi (KI) master field $\Phi_{\pm}$ in (56). However, for $n=2$, the relation between the near-boundary expansions of $\Phi_{\pm}$
@@
is such that the leading and subleading coefficients of $\Phi_{\pm}$ both contain a mixture of sources and vevs for the boundary operators. Explicitly Edalati:2010pn
@@
As shown in Edalati:2010pn, the stress-tensor and current correlators are instead linearly related to a different set of master variables, given by
@@
Since $V_{\pm}(z)$ is not always negative, the potential (166) will generically have singularities at finite $z$, implying that the correlators $\Pi^{\pm}$ associated with $\Psi_{\pm}$ may admit zeroes. Note that, from the definition (165), one can infer that the $\Pi^{\pm}$ are related to the $G^{\pm}$ as Edalati:2010pn
@@ Appendix G Poles of the stress-tensor correlators
We present in this appendix the results of a numerical calculation of the low-lying poles of the helicity-0 stress-tensor correlators analyzed in section 3.2, for $n=3$ boundary spatial dimensions. The caculation was based on the Mathematica package of Jansen:2017oag. As discussed in section 3.2, the poles split into two families: the sound channel associated with the $\Phi^{+}$ master variable, and the diffusive channel associated with $\Phi^{-}$.
@@ G.1 Diffusive channel
Figure 8 shows the low-lying poles of the diffusive channel (which are purely imaginary), for $r_{e}T=(50\pi)^{-1}\simeq 6.4\times 10^{-3}$. The behavior is identical to what was found in Davison:2022vqh for $n=2$: there is a single gapless pole which is diffusive $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, and, up to $\mathcal{O}(r_{e}T)$ corrections, the soft poles equal the IR poles (orange lines)
@@ G.2 Sound channel
up to $\mathcal{O}(r_{e}T)$ corrections. Figure 10 also shows that the sound poles transition from hydrodynamic to relativistic behavior at large momentum, as has been observed previously for $n=2$ Edalati:2010pn.
@@ Appendix H Scalar poles of the magnetic black brane
We present in this appendix the results of a numerical calculation for the low-lying poles of the scalar correlator on the near-extremal magnetic black-brane background discussed in section 3.3. In this case the background itself is numerical, for which the package of Jansen:2017oag does not seem to handle higher working precision, which is necessary to compute more than the first two poles. Our calculation instead used a Mathematica script which is based on the same standard algorithm as Jansen:2017oag, but allows to control the working precision throughout.

### v2

@@ 4 An application to holographic neutrino transport
In this last section, the extremal hydrodynamic product formula (19) is applied to the setup considered in Järvinen et al. (2023). There, a toy model was introduced to investigate neutrino transport in holography, where the charged current spectral functions were computed from the Maxwell equations on the RN background, i.e. the problem considered in section 3.1. The numerical data for the spectral functions was then used to compute the neutrino and anti-neutrino opacities, $\kappa$ and $\bar{\kappa}$. An interesting observation from this calculation was that the opacities at low temperatures were quite well approximated by substituting for the current correlators the leading order hydrodynamic expression Kovtun (2012), extrapolated to zero temperature.
@@
Figure 3: Neutrino opacity $\kappa$ and its hydrodynamic approximations as a function of the neutrino energy $E_{\nu}$, for $\mu/T\simeq 4.7$ (left) and $\mu/T\simeq 65$ (right). $\kappa_{e,0}$ is a normalization factor that was introduced in Järvinen et al. (2023) and $\mu_{\nu}$ the neutrino chemical potential, whose location is indicated by the gray dashed lines.
@@
The better description of the correlators at low energies translates into a better approximation to the neutrino opacities, as illustrated in figures 3 and 4. The latter show a comparison of the exact holographic opacities with the hydrodynamic and extremal hydrodynamic approximations, as a function of neutrino energy $E_{\nu}$ and for two different values of the ratio of chemical potential to temperature [^17] For $T=10\,\text{MeV}$, these values correspond to baryon number densities $n_{B,1}=10^{-3\,}\mathrm{fm^{-3}}$ and $n_{B,2}=1\,\mathrm{fm^{-3}}$, which covers the typical range expected in neutron stars Järvinen et al. (2023).: $\mu_{1}/T\simeq 4.7$ and $\mu_{2}/T\simeq 65$. For most values of $E_{\nu}$, the extremal and standard hydrodynamic approximations are observed to have similar accuracies, except near the dip in the neutrino opacity around $E_{\nu}\sim\mu_{\nu}$ (with $\mu_{\nu}$ the neutrino chemical potential; see figure 3), where the extremal hydrodynamic approximation gives a significantly better description. This is consistent with expectations, since the integrand from which the neutrino opacity is computed (see Järvinen et al. (2023)) is concentrated near $\omega=0$ in this case. Note that this regime of neutrino energies is particularly important to describe well since it corresponds to the effective Fermi surface [^18] See Järvinen et al. (2023) for a discussion of the quasi-particle approximation considered for the description of neutrinos., near which most neutrinos are emitted and absorbed.
@@ 5 Discussion
In this work, we used the holographic product formula introduced in Dodelson et al. (2024) to get insight into the structure of holographic spectral functions in the near-extremal hydrodynamic regime, $\omega,k,T\ll\mu$. In the limit of vanishing temperature $T\ll\omega,k\ll\mu$, the formula provides an expression where both the low temperature gapless modes and the IR scaling behavior appear explicitly, moreover in a simple factorized form. As we showed in sections 3 and 4, the product formula applies in many holographic settings, where it allows to extend the near-extremal hydrodynamic approximation of correlators to the regime of low energies.
@@
  • Another interesting point is related to the recent results of Grozdanov and Vrbica (2025a, b) in four bulk dimensions, which showed that duality relations between different parity sectors of correlators translate into additional constraints on their structure. These may allow to write even more useful expressions in the near-extremal hydrodynamic regime of $(2+1)$-dimensional holographic theories;
  • In Dodelson et al. (2024), the UV limit $\omega\gg T$ of the product formula was analyzed in relation with the OPE in the UV CFT. Here, we were rather interested in an IR regime of the correlators $\omega\ll\mu$, but also considered the limit $T\ll\omega\ll\mu$ to write the extremal hydrodynamic product formula. In this case, the corrections of order $\mathcal{O}(T/\omega)$ to the extremal result may still be controlled by an OPE, although rather associated with the IR CFT, which could be worth investigating;
@@
Finally, the main problem remains to understand what is the structure of the low energy effective field theory (EFT) that underlies the near-extremal spectral functions, which should combine both hydrodynamic-like features and the IR CFT. The formalism of Crossley et al. (2017); Glorioso et al. (2017); Liu and Glorioso (2018) should a priori encompass this kind of theories, since it is well adapted to describe systems with hydrodynamic-like features Blake et al. (2018). This type of effective theories are also known to arise in holographic calculations Nickel and Son (2011); Glorioso et al. (2018); Davison et al. (2023), which however still remain to be extended to the relevant systems.
@@ Appendix A IR limit of near-extremal correlators
Now, the procedure to exhibit the relation with the IR correlator is based on the standard division of the bulk into inner and outer region Faulkner et al. (2011); D’Hoker et al. (2011):
@@
    where the coefficient of the subleading term is the IR correlator by definition Faulkner et al. (2011), and the dots go to zero as $\zeta\to 0$. Note that $\Delta_{\text{IR}}=1/2+\nu$ is (half) the scaling dimension of the IR correlator associated with $\psi$.
  • The outer region is where $\zeta\ll 1$. In particular, it overlaps with the inner region near the boundary of the IR AdSp, where the outer solution behaves as in (127). By linearity of (121), the full outer solution may then be written as Faulkner et al. (2011)
@@ Appendix B An example at next-to-leading order in the near-extremal hydrodynamic expansion
We discuss here how the correlator computed in Goutéraux et al. (2025) at next-to-leading order in the near-extremal hydrodynamic expansion fits with the product formula (14), and in particular with its extremal limit (19).
@@
the result of Goutéraux et al. (2025) takes the form [^22] Note that we also need $i\omega/(2\pi T)\notin\mathbb{N}^{*}$ for $\mathcal{G}$ in (135) to be finite. In other words, the expansion (134) breaks down when $i\omega/(2\pi T)$ is a positive integer.
@@
We are interested in the limit [^23] Note that, in this case, the soft poles do not match the poles of the IR CFT correlator as in (20) Goutéraux et al. (2025). There is therefore no simple expression of the form (23) for the correlator in the general near-extremal regime $|\omega|\sim|\vec{k}^{2}|\sim T\ll r_{e}^{-1}$. However, the extremal formula (19) still applies. $\omega/T\to\infty$, for which the product formula (14) simplifies according to (18), whereas (134)-(135) becomes (for $\omega$ off the negative imaginary axis)
@@
At this order equation (136) has two poles Goutéraux et al. (2025)
@@
with the numerator given by
@@
with the IR scaling dimension given in this case by Goutéraux et al. (2025)
@@ Appendix C Correlators for different field variables in the probe longitudinal gauge sector
For $d\geq 7$, the leading term $\phi_{(0)}$ is non-normalizable and should be indentified as the source, so that $G_{\Psi}^{R}$ is unambiguously identified as $G_{\Psi}^{R}=\phi_{(\Delta)}/\phi_{(0)}$. However, if $d\leq 6$, the dimension $\Delta$ is inferior to $(d+2)/2$, for which two possible quantizations are possible, depending on whether $\phi_{(0)}$ or $\phi_{(\Delta)}$ is treated as the source Klebanov and Witten (1999). In this case, two zero-free product formulae may therefore be written: one for the two-sided correlator associated with $\phi_{(\Delta)}/\phi_{(0)}$, and the other for the inverse $\phi_{(0)}/\phi_{(\Delta)}$ Klebanov and Witten (1999). However, it is not difficult to see that one of these formulae automatically implies the other.
@@ Appendix D Poles of the probe current correlators
In this appendix, we present a numerical calculation of the low-lying (gapless and soft) poles of the RN probe current correlators discussed in section 3.1, in the near-extremal regime. We used for this the Mathematica package of Jansen (2017). The transverse and longitudinal sectors are discussed separately.
@@ D.2 Longitudinal correlator
We now present the numerical results for the poles of the longitudinal correlator, which are shown in figures 6 (imaginary part) and 7 (real part), still with $d=4$ and $r_{e}T=(50\pi)^{-1}$. This sector contains both a diffusive pole $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ and a sequence of soft poles, which interact non-trivially. The behavior as momentum increases is very similar to Goutéraux et al. (2025), with the diffusive pole splitting into two poles with finite real parts before recombining on the imaginary axis several times as it collides with soft poles, before eventually remaining outside the imaginary axis, with a behavior of the form (138). Figure 6 indicates that the soft poles asymptote to the IR poles (153) (orange lines) at large momenta, but deviate significantly around the point where they cross the diffusive mode.
@@ F.1 $n=2$
The helicity-0 correlators on the AdS4-RN background were studied in Edalati et al. (2010a). As for $n=3$, the system of fluctuation equations can be decoupled by considering the Kodama-Ishibashi (KI) master field $\Phi_{\pm}$ in (56). However, for $n=2$, the relation between the near-boundary expansions of $\Phi_{\pm}$
@@
is such that the leading and subleading coefficients of $\Phi_{\pm}$ both contain a mixture of sources and vevs for the boundary operators. Explicitly Edalati et al. (2010a)
@@
As shown in Edalati et al. (2010a), the stress-tensor and current correlators are instead linearly related to a different set of master variables, given by
@@
Since $V_{\pm}(z)$ is not always negative, the potential (166) will generically have singularities at finite $z$, implying that the correlators $\Pi^{\pm}$ associated with $\Psi_{\pm}$ may admit zeroes. Note that, from the definition (165), one can infer that the $\Pi^{\pm}$ are related to the $G^{\pm}$ as Edalati et al. (2010a)
@@ Appendix G Poles of the stress-tensor correlators
We present in this appendix the results of a numerical calculation of the low-lying poles of the helicity-0 stress-tensor correlators analyzed in section 3.2, for $n=3$ boundary spatial dimensions. The caculation was based on the Mathematica package of Jansen (2017). As discussed in section 3.2, the poles split into two families: the sound channel associated with the $\Phi^{+}$ master variable, and the diffusive channel associated with $\Phi^{-}$.
@@ G.1 Diffusive channel
Figure 8 shows the low-lying poles of the diffusive channel (which are purely imaginary), for $r_{e}T=(50\pi)^{-1}\simeq 6.4\times 10^{-3}$. The behavior is identical to what was found in Davison et al. (2023) for $n=2$: there is a single gapless pole which is diffusive $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, and, up to $\mathcal{O}(r_{e}T)$ corrections, the soft poles equal the IR poles (orange lines)
@@ G.2 Sound channel
up to $\mathcal{O}(r_{e}T)$ corrections. Figure 10 also shows that the sound poles transition from hydrodynamic to relativistic behavior at large momentum, as has been observed previously for $n=2$ Edalati et al. (2010a).
@@ Appendix H Scalar poles of the magnetic black brane
We present in this appendix the results of a numerical calculation for the low-lying poles of the scalar correlator on the near-extremal magnetic black-brane background discussed in section 3.3. In this case the background itself is numerical, for which the package of Jansen (2017) does not seem to handle higher working precision, which is necessary to compute more than the first two poles. Our calculation instead used a Mathematica script which is based on the same standard algorithm as Jansen (2017), but allows to control the working precision throughout.